OpenStudy (anonymous):

Simplify: (Ö 3 + Ö -7 )( Ö 3 - Ö -7 )

OpenStudy (anonymous):

5xÖ72 + Ö8x² Break 72 into prime factors 72 = 8·9 = 2·4·3·3 = 2·2·2·3·3 Break 8x² into prime factors 8x² = 2·4·x·x = 2·2·2·x·x Substitute under radicals: _________ _________ 5xÖ2·2·2·3·3 + Ö2·2·2·x·x Group like factors into pairs: _____________ _____________ 5xÖ(2·2)·2·(3·3) + Ö(2·2)·2·(x·x) Write each pair as a square: _______ _______ 5xÖ2²·2·3² + Ö2²·2·x²